Main INDEX
Monthly INDEX
PREV
NEXT
User name E. Cisbani
Log entry time 08:10:12 on June24,2005
Entry number 147075
keyword=Owl Shift Summary
Owl Shift Summary 06/24
01:57-02:05 Problem starting new run 4853, ROC12 did non responded; tryed again with run 4854, same problem. ROC12 rebooted and CODA restarted.
02:57-03:05 Still problem starting new run (4856, 4857), ROC16 did not responded; ROC16 rebooted and CODE restarted.
04:26 Right Septum Magnet tripped due to over-temperature; run 4859 stopped and asked MCC to turn off beam.
04:58 Right Septum Up again
05:20 (around) O2 excess in RICH (see e-log entry 147072)
NOTE: Average Energy Spread (Model) at about 3.8 for the whole shift
Runs Summary:
23:52-00:33 Run 4850 (H2O speed = 599.5)
00:34-01:16 Run 4851 (H2O speed = 605.4)
01:17-01:56 Run 4852 (H2O speed = 602.8), part of run maybe affected by the DAQ problem of the following two runs.
01:57-02:00 Run 4853 and 4854 are junk (DAQ problem)
02:05-02:57 Run 4855 (H2O speed = 602.9), first run after ROC12 reboot and CODA cold restart
02:58-03:01 Run 4856 and 4857 are junk (DAQ problem)
03:05-03:46 Run 4858 (H2O speed = 605.5), first run after ROC16 reboot and CODA cold restart
03:47-04:27 Run 4859 (H2O speed = 603.7), right septum magnet trip
04:59-05:40 Run 4860 (H2O speed = 603.0), first run after right septum recovered
05:41-06:24 Run 4861 (H2O speed = 605.8)
06:25-07:06 Run 4862 (H2O speed = 603.8)
07:07-07:50 Run 4863 (H2O speed = 606.8)
07:52-next shift Run 4864 (H2O speed = 604.4)